# Leveraging Discrete Choice Experiments for User-Centric Requirements Prioritization in mHealth Applications

### v1

@@ 3.1 Key attributes identification
In a DCE, attributes are key features of a product, service, or system presented to participants as variables influencing their choices, and is typically assigned multiple levels, which define the variations of that attribute [klojgaard2012designing]. The attributes are extracted from two primary sources: (i) a SLR concerning the current state of the art of AUIs in mHealth applications [wang2024adaptive], and (ii) the findings of a socio-technical grounded theory analysis that scrutinized data derived from focus groups and interview sessions, focused on examining perceptions regarding the implementation of AUIs in mHealth applications [wang2024adaptiveicse]. The first author identified seven potential attributes from the sources, selecting the six most significant for the DCE survey, which asks participants to assess their importance.
@@ 3.2 Survey development
To study how the attributes in Section 3.1 affect user preferences, we used a factorial design to systematically integrate these levels and evaluate their impact on decision-making [hensher2015applied]. The factorial design comprising six attributes, in which four attributes possess two levels each and two attributes present three levels, results in a full factorial design encompassing 144 ($2^{4}\times 3^{2}=144$) potential choice sets [hensher2015applied]. However, it is impractical to ask participants to make 72 discrete choices due to time and complexity constraints. Thus, we followed established practice and employed a fractional factorial design, reducing the survey to a subset of choice sets [soekhai2019discrete, bridges2011conjoint, weber2021step]. A fractional factorial design selects a smaller subset of combinations from the full factorial set, effectively capturing key effects. Based on D-optimality criteria, our resulting fractional factorial design consists of 18 choice sets [aizaki2008design], converted to nine questions, each offering two choice sets from which respondents need to select. We randomized both the order of the attributes and the sequence of questions presented to participants upon entering the survey to minimize order effects like consistently prioritizing earlier attributes or questions [bridges2011conjoint].
To assess the quality and internal validity of the DCE survey [johnson2019internal], we added a randomly generated tenth question, along with its set of choices, which mirrored the previous question as one attention check question. This facilitates the evaluation of whether the responses exhibited consistency, demonstrated by the selection of the same adaptation design in both equivalent questions. While internal validity checks assess data quality, inconsistent responses to repeated questions align with random utility theory [hess2018revisiting]. To prevent bias or lower statistical efficiency, we do not remove participants failing the validity check, as advised by lancsar2006deleting. We provide information on internal validity to allow the reader to assess possible biases, but responses to the random tenth question are excluded from the analysis [szinay2021understanding]. Figure 2 presents a sample question from the DCE survey, asking participants to choose one of the two presented combined choice sets. We used statistical language R  [^1] https://www.r-project.org and Qualtrics  [^2] https://www.qualtrics.com to design the DCE survey.
@@ 3.3 Piloting and survey adjustments
The pilot study of our DCE survey involved four participants: two experts specializing in statistics and choice modeling, an experienced academic staff member in the field of survey research, and an individual representative of the target demographic, notably possessing both hypertension and lung cancer. The participants filled out the survey, expressed their opinions, and recorded the time it took them to finish, confirming it was within the 20-minute limit. All participants found the DCE survey engaging and relevant but expressed confusion about unclear terminology and how to respond. To address these issues, we made the following revisions: (1) simplified the survey language by using clearer, more intuitive terminology; (2) added a comprehension check questions to ensure participants fully understood each attribute, addressing concerns about potential ambiguity; (3) improved survey instructions by including example questions to better guide participants through the response process and (4) added one more attention check question to ensure thoughtful responses. The survey consists of three sections, guiding participants through the study step by step. Section A begins with an introduction outlining the study’s purpose. It also includes a screening step to confirm participant eligibility (i.e., experience with mHealth apps and a chronic condition), followed by demographic questions such as age, gender, and other relevant background details. Section B introduces participants to the concept of AUI in mHealth applications, explaining their relevance to the study. This section also provides an overview of the DCE, including a sample question and a description of the attributes under investigation. To ensure participants understand these concepts before proceeding, a comprehension check is included at the end of this section, allowing up to three attempts. Participants then complete ten DCE choice tasks, with the tenth question serving as an attention check to assess internal response consistency. Section C concludes with five questions about participants’ health conditions and their attitudes toward managing health challenges, along with a second attention check to ensure continued engagement with the survey content. A complete version of the survey instrument is available in the supplementary materials for reference [^3] https://doi.org/10.5281/zenodo.17555183. In total, the survey includes three quality checks: one comprehension check (Section B), one attention check within the DCE (Section B), and one final attention check (Section C).
@@ 3.4 Data collection
The data collection process for our DCE survey was conducted in two stages. Stage One involved non-probability sampling through community and organizational outreach within Australia. The survey was distributed online via social media platforms and disseminated through national and state-level non-governmental health organizations, including the Stroke Foundation, Victoria Diabetes, Dementia Australia, and Kidney Health Australia. In addition, physical posters were displayed in private clinics and select hospitals. Collaboration with JOIN US  [^4] https://www.joinus.org.au facilitated the promotion of the study to individuals with specific health conditions. A total of 122 responses were collected through these channels during stage one.
In stage two, we employed a purposive sampling strategy through Prolific [^5] https://www.prolific.com/ by employing a custom screener to refine the pool of participants to individuals with chronic diseases. This change was essential for increase the sample size and improved diversity in participant demographics since all stage one participants were from Australia. We excluded participants with discrepancies between their screening and prescreening responses on Prolific to identify falsified answers.A total of 64 responses were collected through Prolific during stage two. Overall, we collected 186 responses. To determine the required sample size for a DCE, orme1998sample proposes a commonly used rule of thumb estimating a minimum of 94 respondents for our survey design, while research suggests that response precision improves significantly with around 150 participants [johnson2013constructing], despite the challenges in calculating precise sample sizes for DCEs. Based on this guideline, we set 150 as our benchmark and successfully exceeded this target, ensuring robust and reliable findings.
@@ 3.5 Data analysis
Individual characteristics were explored descriptively. To analyze the data collected from the DCE, we used statistical models designed to understand how participants make trade-offs between different features. These models help quantify the importance of each attribute by examining the choices participants made across multiple scenarios. A mixed logit model was used as this modeling approach accommodates multiple observations per participant and does not require an assumption of independence of irrelevant alternatives [greene2003latent]. The analysis was performed in R  [^6] https://www.r-project.org, specifically using the mlogit package based on 1000 Halton draws to ensure stable coefficient estimates. A univariate regression analysis is conducted to examine the independent effect of the participant’s characteristics on choices, which helps to identify the significant factors that affect preference. Following this, a multivariate regression model is fitted, incorporating all attributes and participant characteristic factors identified in the univariate analysis, to evaluate the joint effects of multiple factors while controlling for confounding and capturing the relative importance of each.
Subsequently, an analysis specific to subgroups was conducted to facilitate a direct comparison of preferences across predefined cohorts. Standardized Relative Attribute Importance (RAI) was used to elucidate the relative contribution of various attributes in distinct subgroups. A Marginal Rate of Substitution (MRS) analysis was conducted to evaluate the trade-offs between attributes by estimating a mixed logit model to derive attribute coefficients. The usability of the app is chosen as a reference attribute, given its high significance and its crucial role in the study. Subgroup analyzes are performed to capture heterogeneity in preferences between groups, with separate models estimated for each subgroup. MRS values are normalized, ensuring that the reference attribute had a relative utility of 100% for ease of comparison. The study examined the influence of seven predetermined participant characteristics on response data. These characteristics were: nationality (Australia or Non-Australia); age; gender; educational attainment (below Bachelor’s degree or above Bachelor’s degree); health status (Poor (including fair) or Good); and stress coping mechanism (positive (including mixed feelings) or negative).
@@ 4 Results
4.1 Participants
@@
Table 2 presents the demographic and key characteristics of our DCE study’s 186 participants, who were predominantly women (60%) and mainly residing in Australia (66%). Our participants ranged in age, with the largest group aged 65–74 (24%), followed by those aged 35–44 (20%) and 25–34 (19%). The majority identified as White (82%), with smaller proportions identifying as Multicultural (7%), Black or African American (6%), Asian (3%) or Aboriginal / Torres Strait Islander (1%). Most of the participants (56%) did not have a bachelor’s degree, while 27% had a bachelor’s degree and 16% had advanced degrees. Regarding health, 41% rated their health as fair, while 32% rated it good, and the remaining categories ranged from poor (16%) to very poor (5%). With 62% of individuals reporting cardiometabolic and endocrine disorders, respiratory conditions (29%), and neurological or mental health issues (22%). A mix of positive thinking and avoidance is the most common (33%), followed by staying positive and handling stress (31%). This diverse sample provides broad information on the demographic and health characteristics of the population.
@@ 4.2 Establishing initial preferences
Table 3: Main results of mixed logit model
@@
In the discrete choice analysis, all attributes are statistically different from 0, indicating their importance in shaping user preferences for adaptation design. Table 3 presents the estimated coefficients, standard errors, z-values, 95% confidence intervals (CIs), and standard deviations (SDs) of the random parameters from the mixed logit model. The mixed logit model was estimated using maximum simulated likelihood with 1,000 Halton draws. The final log-likelihood was –1051.5, with a McFadden’s pseudo-R² of 0.113, indicating a moderate model fit. The AIC and BIC values were 2119.865 and 2184.946, respectively. Standard deviations of the random coefficients reveal considerable heterogeneity in preferences across individuals. This suggests that participants vary substantially in how they value these features, justifying the use of a random-parameter approach to better capture individual-level variation. The regression coefficients for each attribute level represent the mean part-worth utility of that attribute level in the respondent sample. A positive value denotes utility/satisfaction, and a negative value denotes disutility/dissatisfaction. The usability of the app is the most important factor influencing user preferences ($\beta=0.704$, 95% CI 0.579–0.829), highlighting the critical role of maintaining usability while adapting the UI. This is followed by controllability over adaptations ($\beta=0.279$, 95% CI 0.085–0.473), frequency of adaptations ($\beta=0.263$, 95% CI 0.092–0.435), and granularity of adaptations ($\beta=0.221$, 95% CI 0.021–0.421), which reflect the value users place on the flexibility of adaptation design. In contrast, the involvement of caregivers ($\beta=-0.33$, 95% CI 0.489-0.489 to 0.171-0.171) and patterns of function usage ($\beta=-0.229$, 95% CI 0.383-0.383 to 0.074-0.074) posed significant obstacles to the adoption of the adaptation design. This suggests that users may be averse to adaptations occurring in features they frequently use, while the involvement of caregivers may lead to a perceived redundancy in adaptations, as users can obtain assistance in utilizing the application from their caregivers. In our DCE, due to the innovative nature of research concerning the adoption of adaptive designs in mHealth technology and the absence of empirical evidence indicating possible attribute interactions, we opted to focus solely on the main effects.
@@
Figure 3: Radar chart for subgroup comparison: This radar chart illustrates the Relative Attribute Importance (RAI) for three key attributes—controllability over adaptations, usability of the app, and granularity of adaptations—across subgroups based on gender, health condition, age, and coping mechanisms. Lines represent the relative weight each subgroup placed on each attribute. Values are point estimates for descriptive comparison only; no statistical tests were conducted.
@@ 4.3 Trade off between different attributes
To assess how much participants were willing to sacrifice usability for enhancements in other features, we measured this trade-off using the MRS relative to usability. Table 4 displays the MRS values, where a higher positive MRS indicates a greater willingness to trade usability for that attribute, and a negative MRS suggests the attribute is viewed as less valuable or undesirable compared to usability. We found that participants are highly willing to trade usability for controllability over adaptations (40%) and frequency of adaptations (37%), highlighting their importance. The involvement of caregivers (-5%) and function usage patterns (-0.3%) are perceived unfavorably, indicating that participants regard these factors as less desirable or potentially counterproductive in terms of usability. However, the slight negative valuation of these attributes indicates that they hold a relative importance that is closer to usability.
@@
We also performed a subgroup analysis by age, gender, health condition, and coping mechanism based on our previous analysis in Section 4.2.1. Subgroup MRS values were obtained by normalizing point estimates from subgroup-specific models relative to the usability attribute. Confidence intervals for MRS values were not computed, and no formal statistical tests were conducted; therefore, all subgroup comparisons should be interpreted as descriptive rather than inferential. Women prioritized infrequent adaptations (45%) and controllability (24%), compared to their male counterparts. Male participants would require an increase of 10% in usability to compensate for an increase in caregiver involvement. This suggests that caregiver involvement is viewed as a detractor or a feature that reduces the perceived value of the system for men. Young participants value more controllability compared to their older counterparts (54% versus 25%). In contrast, older participants prioritize infrequent adaptations (49%) and small-scale adaptations (35%). Those in good health prioritize controllability over adaptations (57%) more than those in poor health (33%), reflecting a stronger desire for autonomy. Participants in poor health place greater importance on the granularity of adaptations (36% versus 22%), emphasizing the need for precision and small-scale changes, while also valuing caregiver involvement (12%) and less frequent adaptations (4%). Different approaches to coping led to varied preferences. Those who used positive coping strategies placed a high priority on controllability (85%), whereas they assigned less significance to infrequent change (28%). In contrast, participants with negative coping strategies preferred infrequent change (44%) and smaller changes (36%) over controllability (19%). Participants with a positive coping attitude are more open to compromises, allowing caregivers to utilize the app (19%), whereas those with a negative coping style tend to ignore their health conditions and resist involving the caregiver in the app’s usage (-17%).
@@
Controllability over adaptations ($40\%$ ($24\%$ ($55\%$ ($54\%$ ($25\%$ ($57\%$ ($33\%$ ($85\%$ ($19\%$
Involvement of caregivers -$5\%$ -$\% -0.10.1\%$ -$6\%$ -$3\%$ -$14\%$ (1212% ($19\%$ -$17\%$
Function usage patterns -$0.3\%$ ($1\%$ -$2\%$ -$1.8\%$ ($1.8\%$ -$1.2\%$ ($4\%$ ($0.2\%$ -$1\%$
Frequency of adaptations ($37\%$ ($31\%$ ($45\%$ ($26\%$ ($49\%$ ($41\%$ ($39\%$ ($28\%$ ($44\%$
Granularity of adaptations $\% (3131\%$ ($32\%$ ($28\%$ ($35\%$ ($22\%$ ($36\%$ ($22\%$ ($36\%$
a. The table represents the Marginal Rate of Substitution (MRS) expressed as percentages, derived from the mixed logit regression model. The percentages indicate the degree to which participants are willing to trade off usability (the reference attribute) for improvements in other attributes. (  Positive values: Indicate that participants are willing to compromise usability to prioritize improvements in that attribute. Higher percentages reflect greater willingness to trade usability for improvements in the respective attribute. ”  Negative values: Indicate that participants are willing to sacrifice that attribute to preserve usability, implying lower desirability or priority of the attribute relative to usability.

### v2

@@ 3.1 Key attributes identification
In a DCE, attributes are key features of a product, service, or system presented to participants as variables influencing their choices, and are typically assigned multiple levels [57]. The attributes are extracted from two primary sources: (i) a SLR concerning the current state of the art of AUIs in mHealth applications [97], and (ii) the findings of a socio-technical grounded theory analysis that scrutinized data derived from focus groups and interview sessions, focused on examining perceptions regarding the implementation of AUIs in mHealth applications [98]. The first author identified seven potential attributes from the sources, selecting the six most significant for the DCE survey, which asks participants to assess their importance.
@@ 3.2 Survey development
To study how the attributes in Section 3.1 affect user preferences, we used a factorial design to systematically integrate these levels and evaluate their impact on decision-making [43]. The factorial design comprising six attributes, in which four attributes possess two levels each and two attributes present three levels, results in a full factorial design encompassing 144 ($2^{4}\times 3^{2}=144$) potential choice sets [43]. However, it is impractical to ask participants to make 72 discrete choices due to time and complexity constraints. Thus, we followed established practice and employed a fractional factorial design, reducing the survey to a subset of choice sets [86, 16, 99]. A fractional factorial design selects a smaller subset of combinations from the full factorial set, effectively capturing key effects. Based on D-optimality criteria, our resulting fractional factorial design consists of 18 choice sets [4], converted to nine questions, each offering two choice sets from which respondents need to select. We randomized both the order of the attributes and the sequence of questions presented to the participants when entering the survey to minimize order effects, such as consistently prioritizing earlier attributes or questions [16].
To assess the quality and internal validity of the DCE survey [51], we added a randomly generated tenth question, along with its set of choices, which mirrored the previous question as one attention check question. This facilitates the evaluation of whether the responses exhibited consistency, demonstrated by the selection of the same adaptation design in both equivalent questions. Although internal validity checks assess data quality, inconsistent responses to repeated questions align with random utility theory [44]. To avoid bias or lower statistical efficiency, we do not remove participants who fail the validity check, as advised by Lancsar and Louviere [59]. We provide information on internal validity to allow the reader to assess possible biases, but responses to the random tenth question are excluded from the analysis [89]. Figure 2 presents a sample question from the DCE survey, asking participants to choose one of the two presented combined choice sets. We used statistical language R  [^1] https://www.r-project.org and Qualtrics  [^2] https://www.qualtrics.com to design the DCE survey.
@@ 3.3 Piloting and survey adjustments
The pilot study of our DCE survey involved four participants: two experts specializing in statistics and choice modeling, an experienced academic staff member in the field of survey research, and an individual representative of the target demographic, notably possessing both hypertension and lung cancer. The participants filled out the survey, expressed their opinions, and recorded the time it took them to finish, confirming it was within the 20-minute limit. All participants found the DCE survey engaging and relevant but expressed confusion about unclear terminology and how to respond. To address these issues, we made the following revisions: (1) simplified the survey language by using clearer, more intuitive terminology; (2) added one comprehension check questions to ensure participants fully understood each attribute, addressing concerns about potential ambiguity; (3) improved survey instructions by including example questions to better guide participants through the response process; and (4) added one more attention check question to ensure thoughtful responses.
The survey consists of three sections that guide the participants through the study step by step. Section A begins with an introduction explaining the purpose of the study. It also includes a screening step to confirm participant eligibility (i.e., experience with mHealth apps and a chronic disease), followed by demographic questions such as age, gender, and other relevant background details. Section B introduces participants to the concept of AUI in mHealth applications, explaining their relevance to the study. This section also provides an overview of the DCE, including a sample question and a description of the attributes under investigation. To ensure that participants understand these concepts before proceeding, a comprehension check is included at the end of this section, allowing up to three attempts. The participants then complete ten DCE choice tasks, with the tenth question serving as an attention check to assess the consistency of the internal response. Section C concludes with five questions on participants’ health conditions and their attitudes towards managing health challenges, along with a second attention check to ensure continued engagement with the survey content. A complete version of the survey instrument is available in the supplementary materials for reference [^3] https://doi.org/10.5281/zenodo.17555183. In total, the survey includes three quality checks: one comprehension check (Section B), one attention check within the DCE (Section B), and one final attention check (Section C).
@@ 3.4 Data collection
The data collection process for our DCE survey was conducted in two stages. Stage One involved non-probability sampling through community and organizational outreach within Australia. The survey was distributed online through social media platforms and distributed through national and state-level non-governmental health organizations, including the Stroke Foundation, Victoria Diabetes, Dementia Australia and Kidney Health Australia. In addition, physical posters were displayed in private clinics and select hospitals. Collaboration with JOIN US  [^4] https://www.joinus.org.au facilitated the promotion of the study to individuals with specific health conditions. A total of 122 responses were collected through these channels during stage one.
In stage two, we employed a purposive sampling strategy through Prolific [^5] https://www.prolific.com/ by employing a custom screener to refine the pool of participants to individuals with chronic diseases. This change was essential to increase the sample size and improve diversity in participant demographics since all stage one participants were from Australia. We excluded participants with discrepancies between their screening and prescreening responses on Prolific to identify falsified answers.A total of 64 responses were collected through Prolific during stage two. Overall, we collected 186 responses. To determine the required sample size for a DCE, Orme [68] proposes a commonly used rule of thumb estimating a minimum of 94 respondents for our survey design, while research suggests that response precision improves significantly with around 150 participants [50], despite the challenges in calculating precise sample sizes for DCEs. Based on this guideline, we set 150 as our benchmark and successfully exceeded this target, ensuring robust and reliable findings.
@@ 3.5 Data analysis
Individual characteristics were explored descriptively. To analyze the data collected from the DCE, we used statistical models designed to understand how participants make trade-offs between different features. These models help quantify the importance of each attribute by examining the choices that participants made in multiple scenarios. A mixed logit model was used as this modeling approach accommodates multiple observations per participant and does not require the assumption of the independence of irrelevant alternatives [34]. The analysis was performed in R  [^6] https://www.r-project.org, specifically using the mlogit package based on 1000 Halton draws to ensure stable coefficient estimates. A univariate regression analysis is conducted to examine the independent effect of the participant’s characteristics on choices, which helps to identify the significant factors that affect preference. Following this, a multivariate regression model is fitted, incorporating all attributes and participant characteristic factors identified in the univariate analysis, to evaluate the joint effects of multiple factors while controlling for confounding and capturing the relative importance of each.
Subsequently, an analysis specific to subgroups was conducted to facilitate a direct comparison of preferences across predefined cohorts. The standardized Relative Attribute Importance (RAI) was used to elucidate the relative contribution of various attributes in distinct subgroups. A Marginal Rate of Substitution (MRS) analysis was conducted to evaluate the trade-offs between attributes by estimating a mixed logit model to derive attribute coefficients. The usability of the app is chosen as a reference attribute, given its high significance and its crucial role in the study. Subgroup analyzes are performed to capture heterogeneity in preferences between groups, with separate models estimated for each subgroup. MRS values are normalized, ensuring that the reference attribute had a relative utility of 100% for ease of comparison. The study examined the influence of seven predetermined participant characteristics on response data. These characteristics were: nationality (Australia or Non-Australia); age; gender; educational attainment (below Bachelor’s degree or above Bachelor’s degree); health status (Poor (including fair) or Good); and stress coping mechanism (positive (including mixed feelings) or negative).
@@ 4 Results
This section presents the findings of the DCE, including the’ overall preferences for adaptive mHealth features, subgroup differences based on demographic and psychological factors, and the trade-offs users are willing to make between competing design attributes.
4.1 Participants demographic
@@
Table 2 presents the demographic and key characteristics of our DCE study’s 186 participants, who were predominantly women (60%) and mainly residing in Australia (66%). Our participants ranged in age, with the largest group aged 65–74 (24%), followed by those aged 35–44 (20%) and 25–34 (19%). The majority identified as White (82%), with smaller proportions identifying as Multicultural (7%), Black or African American (6%), Asian (3%) or Aboriginal / Torres Strait Islander (1%). Most of the participants (56%) did not have a bachelor’s degree, while 27% had a bachelor’s degree and 16% had advanced degrees. Regarding health, 41% rated their health as fair, while 32% rated it good, and the remaining categories ranged from poor (16%) to very poor (5%). With 62% of the people reporting cardiometabolic and endocrine disorders, respiratory conditions (29%) and neurological or mental health problems (22%). A mix of positive thinking and avoidance is the most common (33%), followed by staying positive and managing stress (31%). This diverse sample provides broad information on the demographic and health characteristics of the population.
@@ 4.2 Establishing initial preferences
Table 3: Estimated user preferences for adaptive mHealth features based on mixed logit model
@@
This subsection presents the main effects from the mixed logit model to identify overall user preferences for adaptive mHealth features, followed by subgroup analyses that examine how these preferences vary across health condition, gender and coping mechanism. In the discrete choice analysis, all attributes are statistically different from 0, indicating their importance in shaping user preferences for adaptation design. Table 3 presents the estimated coefficients, standard errors, z-values, 95% confidence intervals (CI) and standard deviations (SD) of the random parameters of the mixed logit model. The mixed logit model was estimated using the maximum simulated likelihood with 1,000 Halton draws. The final log-likelihood was –1051.5, with a McFadden’s pseudo-R² of 0.113, indicating a moderate model fit. The AIC and BIC values were 2119.865 and 2184.946, respectively. The standard deviations of the random coefficients reveal considerable heterogeneity in preferences between individuals. This suggests that participants vary substantially in how they value these features, which justifyes the use of a random-parameter approach to better capture individual-level variation. The regression coefficients for each attribute level represent the mean part-worth utility of that attribute level in the respondent sample. A positive value denotes utility/satisfaction and a negative value denotes disutility/dissatisfaction. The usability of the app is the most important factor influencing user preferences ($\beta=0.704$, 95% CI 0.579–0.829), highlighting the critical role of maintaining usability while adapting the UI. This is followed by controllability over adaptations ($\beta=0.279$, 95% CI 0.085–0.473), frequency of adaptations ($\beta=0.263$, 95% CI 0.092–0.435) and granularity of adaptations ($\beta=0.221$, 95% CI 0.021–0.421), reflecting the value users place on the flexibility of adaptation design. In contrast, the involvement of caregivers ($\beta=-0.33$, 95% CI 0.489-0.489 to 0.171-0.171) and patterns of function usage ($\beta=-0.229$, 95% CI 0.383-0.383 to 0.074-0.074) posed significant obstacles to the adoption of the adaptation design. This suggests that users may be averse to adaptations occurring in features they frequently use, while the involvement of caregivers may lead to a perceived redundancy in adaptations, as users can obtain assistance in utilizing the application from their caregivers. In our DCE, due to the innovative nature of research concerning the adoption of adaptive designs in mHealth technology and the absence of empirical evidence indicating possible attribute interactions, we opted to focus solely on the main effects.
@@
Figure 3: Radar chart for subgroup comparison. This radar chart illustrates the Relative Attribute Importance (RAI) for three key attributes—controllability over adaptations, usability of the app, and granularity of adaptations—across subgroups based on gender, health condition, age, and coping mechanisms. Lines represent the relative weight each subgroup placed on each attribute. Values are point estimates for descriptive comparison only; no statistical tests were conducted.
@@ 4.3 Trade off between different attributes
This subsection examines the trade-offs users are willing to make between usability and other adaptation attributes, using MRS to quantify these preferences. The definition and calculation of MRS are detailed in Section 3.5. Table 4 displays the MRS values, where a higher positive MRS indicates a greater willingness to trade usability for that attribute, and a negative MRS suggests the attribute is viewed as less valuable or undesirable compared to usability. We found that participants are highly willing to trade usability for controllability over adaptations (40%) and frequency of adaptations (37%), highlighting their importance. The involvement of caregivers (-5%) and function usage patterns (-0.3%) are perceived unfavorably, indicating that participants regard these factors as less desirable or potentially counterproductive in terms of usability. However, the slight negative valuation of these attributes indicates that they have a relative importance that is closer to usability.
@@
We also performed a subgroup analysis by age, gender, health condition, and coping mechanism based on our previous analysis in Section 4.2.1. Subgroup MRS values were obtained by normalizing point estimates from subgroup-specific models relative to the usability attribute. The confidence intervals for the MRS values were not computed and no formal statistical tests were conducted; therefore, all subgroup comparisons should be interpreted as descriptive rather than inferential. Women prioritized infrequent adaptations (45%) and controllability (24%), compared to their male counterparts. Male participants would require an increase of 10% in usability to compensate for an increase in caregiver involvement. This suggests that caregiver involvement is viewed as a detractor or a feature that reduces the perceived value of the system for men. Young participants value more controllability compared to their older counterparts (54% versus 25%). In contrast, older participants prioritize infrequent adaptations (49%) and small-scale adaptations (35%). Those in good health prioritize controllability over adaptations (57%) more than those in poor health (33%), reflecting a stronger desire for autonomy. Participants in poor health place greater importance on the granularity of adaptations (36% versus 22%), emphasizing the need for precision and small-scale changes, while also valuing caregiver involvement (12%) and less frequent adaptations (4%). Different approaches to coping led to varied preferences. Those who used positive coping strategies placed a high priority on controllability (85%), whereas they assigned less significance to infrequent change (28%). In contrast, participants with negative coping strategies preferred infrequent change (44%) and smaller changes (36%) over controllability (19%). Participants with a positive coping attitude are more open to compromises, allowing caregivers to utilize the app (19%), whereas those with a negative coping style tend to ignore their health conditions and resist involving the caregiver in the app’s usage (-17%).
@@
Controllability over adaptations $40\%$ $24\%$ $55\%$ $54\%$ $25\%$ $57\%$ $33\%$ $85\%$ $19\%$
Involvement of caregivers -$5\%$ -$\% -0.10.1\%$ -$6\%$ -$3\%$ -$14\%$ 1212% $19\%$ -$17\%$
Function usage patterns -$0.3\%$ $1\%$ -$2\%$ -$\% 1.81.8\%$ -$1.2\%$ $4\%$ $0.2\%$ -$1\%$
Frequency of adaptations $37\%$ $31\%$ $45\%$ $26\%$ $49\%$ $41\%$ $39\%$ $28\%$ $44\%$
Granularity of adaptations $\% 3131\%$ $32\%$ $28\%$ $35\%$ $22\%$ $36\%$ $22\%$ $36\%$
a. The table represents the Marginal Rate of Substitution (MRS) expressed as percentages, derived from the mixed logit regression model. The percentages indicate the degree to which participants are willing to trade off usability (the reference attribute) for improvements in other attributes. Positive values: Indicate that participants are willing to compromise usability to prioritize improvements in that attribute. Higher percentages reflect greater willingness to trade usability for improvements in the respective attribute. Negative values: Indicate that participants are willing to sacrifice that attribute to preserve usability, implying lower desirability or priority of the attribute relative to usability.